Refraction of a plane wave at the interface between two dielectrics.

1
Expert's answer
2022-01-03T14:34:05-0500

We know that


Incidence wave"E=E_0e^{j(k.r-wt)}"

Reflected wave

"E=E_{01}e^{j(k_1.r-w_1t)}"

Refracted wave

"E_2=E_{02}e^{j(k_2.r-w_2t)}"

"k_1" =Reflected wave propagation constant

"k_2=" Refracted wave propagation constant

"k" =Incidence wave Propagation constant

tangential of E


"(E_0)_te^{(k.r-wt)}+(E_0)_te^{(k_1.r-w_1t)}=(E_0)_te^{(k_2.r-w_2t)}"

"k.r=k_1.r=k_2.r\\k_x=k_{1x}=k_{2x}"

For refracted


"k_{2x}=k_{1x}"

"k_{1x}sin\theta_1=k_{1x}sin\theta_2"

"\frac{w_1}{v_1}sin\theta_1=\frac{w_2}{v_2}sin\theta_2"

"w_1=w_2"

"\frac{c}{v_1}sin\theta_1=\frac{c}{v_2}sin\theta_2"

"n_1sin\theta_1=n_2sin\theta_2"

Refracted law